OpenStudy (anonymous):

Dave Horn invested half of his money at 5%,one third of his money at 4%, and the rest of his money at 3.5%.If his total annual investment income was $530,how much had he invested?

OpenStudy (anonymous):

is the nswer 12000.

OpenStudy (anonymous):

answer*

OpenStudy (anonymous):

checks out to me. At 12,000. half of his money is 6,000x0.05=300 1/3 of his money is 4000x0.04=160 12000-6000+4000=12000-10000=2000 2000x0.035=70 300+160+70=530

OpenStudy (anonymous):

:) congrats.